John is drawing a regular polygon. He has drawn two sides with an exterior angle measure of 45°, as shown below.
When John completes his regular polygon, what will be the sum, in degrees, of the measures of the interior angles?

Answer by Alan3354(69443)

John is drawing a regular polygon. He has drawn two sides with an exterior angle measure of 45°, as shown below.
-------
Nothing is shown.
---------
When John completes his regular polygon, what will be the sum, in degrees, of the measures of the interior angles?
---
The sum of exterior angles of ALL polygons is 360 degrees.
360/45 --> 8 sides
Each interior angle = 180 - 45 = 135 degs
8*135 = 1080 degrees